OpenStudy (anonymous):

Is this right? (I'm using L'H rule) lim ln(x)sin(x) = ln(x)/1/sin(x) x->0^+ L'H = 1/x (1(cos(x))/ sin^(2)(x)) = sin^(2)(x)/xcos(x) = lim(x) tan(x)sin(x)/x x->0^(+) = lim (sec^(2)(x)sin(x) + tan(x)cos(x))/1 L'H x->0^(+) = 0

OpenStudy (anonymous):

Note I was lazy with limit notation :L

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0^+} \ln(x)\sin(x)= \lim_{x \rightarrow 0^+} \frac{\ln(x)}{\csc(x)}=\lim_{x \rightarrow 0^+} \frac{\frac{1}{x}}{-\csc(x)\cot(x)}=\lim_{x \rightarrow 0^+}\frac{-\sin^2(x)}{x \cos(x)}\] You are again at a 0/0 so you need to use it again... \[\lim_{x \rightarrow 0^+}\frac{-(\cos(x)\tan(x)+\sin(x)\sec^2(x))}{1}=0\] Sorry for typing it, i work it out as I type it.